INAF LBC queue observing, 2013 January 14-15

Observer: M.Faccini, R. Speziali, F.Cusano
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: M. Edwars
Instrument: LBC

Summary

At sunset temperature is about -17/-18, clouds and few snow.
Weather seems variable as foreseen.
Last night problems with MODS still not solved, at 0:30 UT technicians are still working on LBC-red to solve the problem about winter position of the cameras.
at the beginning of the night the sky is clear and ext temperature is about -19/-20 at less than 1 degree from dewpoint temperature. We can not open the dome.
Around 7 UT temperature increase to -17 and dewpoint is at -25, but we have snow on the roof. The Dome will remain close due to snow on the roof.

-- MarcoFaccini - 15 Jan 2013
Topic revision: r3 - 15 Jan 2013, MarcoFaccini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback